# Amortized Variational Transdimensional Inference

### v1

@@ Appendix A Construction and analysis of a CoSMIC normalizing flow
(a) $\qquad\tilde{q}_{\phi}^{(P_{m})}(\bm{\theta}_{m},\bm{u}_{\setminus m}\mid m)\;=\;q_{\phi}(\bm{\theta}_{m}\mid m)\;\nu_{d_{\setminus m}}\!\bigl(\bm{u}_{\setminus m}\bigr),$
and so (b) $\qquad\int\tilde{q}_{\phi}^{(P_{m})}(\bm{\theta}_{m},\bm{u}_{\setminus m}\mid m)\,d\bm{u}_{\setminus m}=q_{\phi}(\bm{\theta}_{m}\mid m).$
Write $T_{\phi}^{(P_{m}),-1}=(T_{\phi,m}^{-1},\operatorname{Id})$, where $\operatorname{Id}$ denotes the identity transform, and let the permuted reference vector be $P_{m}\bm{z}=(\bm{z}_{m},\bm{z}_{\setminus m})\in\mathbb{R}^{{d_{m}}}\times\mathbb{R}^{d_{\setminus m}}$. Because the masking function $C$ sets every transform $\tau_{\bm{\rho}_{i}}$ with $C_{i}(m)=0$ to the identity, the inverse flow splits as
$T_{\phi}^{(P_{m}),-1}(\bm{\theta}_{m},\bm{u}_{\setminus m})\;=\;\bigl(T_{\phi,m}^{-1}(\bm{\theta}_{m}),\;\bm{u}_{\setminus m}\bigr),$
where $T_{\phi,m}^{-1}:\Theta_{m}\to\mathbb{R}^{{d_{m}}}$ is the active block and the dummy block is exactly the identity. Consequently, the Jacobian matrix of $T_{\phi}^{(P_{m}),-1}$ is block upper‑triangular with $\det\nabla T_{\phi}^{(P_{m}),-1}=\det\nabla T_{\phi,m}^{-1}\times 1.$
@@
$\tilde{q}_{\phi}^{(P_{m})}(\bm{\theta}_{m},\bm{u}_{\setminus m}\!\mid\!m)$ $=\nu_{{d_{\mathrm{max}}}}\!\bigl(T_{\phi}^{(P_{m}),-1}(\bm{\theta}_{m},\bm{u}_{\setminus m})\bigr)\bigl|\det\nabla T_{\phi}^{(P_{m}),-1}\bigr|$
@@
$=D_{\mathrm{KL}}(q_{\psi,\phi}||\pi)$
$=\pazocal{L}(\psi,\phi).$
A.1 CoSMIC IAF computational complexity
Corollary A.1 (Computational complexity).
@@ A.1 CoSMIC IAF computational complexity
A.2 Experimental CoSMIC transform compositions
The experiments use the below compositions of transforms as inverse autoregressive flows $T_{\phi}(\bm{z}\mid m)$ where $\bm{z}$ are the inputs from the reference distribution and $m$ is the context input. All compositions except for the diagonal Gaussian are assumed to have the strict left-right permutations discussed in Appendix˜A. The term “block” is defined in Section˜A.3.
Context encoder:
Model-specific reverse-permutation:
Flow compositions commonly include reverse permutations to ensure expressibility of an autoregressive-NN-based flow is (approximately) the same for all coordinates. Denoting the generic reverse permutation for all coordinates as $P^{\operatorname{rev}}$, we assume the strict left-right permutation $P_{m}$ (as per Appendix˜A) has been applied, and hence define the left-most $d_{m}$-coordinate reverse permutation $P_{<d_{m}}^{\operatorname{rev}}$.
Affine(5,5):
The learned component is the affine masked autoregressive transform (Papamakarios et al., , 2017), denoted here as $T_{\phi_{k}}^{\operatorname{Affine}}$ for transforms $k=1,\dots,5$, each having 5 blocks. We set $T_{\phi}:=T_{\phi_{5}}^{\operatorname{Affine}}\circ P_{<d_{m}}^{\operatorname{rev}}\circ\dots\circ P_{<d_{m}}^{\operatorname{rev}}\circ T_{\phi_{1}}^{\operatorname{Affine}}$.
Spline(4,6):
The learned component is the rational quadratic spline masked autoregressive flow architecture (Durkan et al., , 2019), denoted here as $T_{\phi_{k}}^{\operatorname{RQ-Spline}}$. Each $T_{\phi_{k}}^{\operatorname{RQ-Spline}}$ has 6 blocks. Additionally, we define a fixed global affine transform $T_{\mu_{g},\sigma_{g}}$ that is not dependent on inputs nor context and hence has only two learnable parameters: scale $\mu_{g}$ and shift $\sigma_{g}$. We set $T_{\phi}:=T_{\mu_{g},\sigma_{g}}\circ T_{\phi_{5}}^{\operatorname{RQ-Spline}}\circ P_{<d_{m}}^{\operatorname{rev}}\circ\dots\circ P_{<d_{m}}^{\operatorname{rev}}\circ T_{\phi_{1}}^{\operatorname{RQ-Spline}}$.
A.3 Autoregressive flow definitions
We use the residual variant of the Masked Autoencoder for Distribution Estimation (MADE) (Germain et al., , 2015), implemented in PyTorch by (Durkan et al., , 2020). Each block maintains the autoregressive property by assigning degrees $\{1,\dots,d\}$ to inputs and propagating them forward.
Appendix B Theoretical analysis of the model weights distribution
@@ B.1 Regularity assumptions
This mild assumption is satisfied, for example, when $\nu$ is a zero-mean Gaussian distribution and $\tilde{f}$ is Lipschitz continuous on its first argument, in which case $\sigma_{\epsilon}$ only depends on $\tilde{f}$ through its Lipschitz constant (Boucheron et al., , 2013; Pisier, , 2016).
B.2 Gaussian process model
@@
B.3 Upper confidence bound (UCB) algorithm
Given the GP posterior, we formulate an upper confidence bound algorithm (Srinivas et al., , 2010) with:
@@ B.3 Upper confidence bound (UCB) algorithm
B.4 Approximation errors under sub-Gaussian noise
@@
where $\kappa(x,\pazocal{X}_{n}):=[\kappa(x,x_{1}),\dots,\kappa(x,x_{n})]$, $\mathbf{K}_{n}:=[\kappa(x_{i},x_{j})]_{i,j=1}^{n}$, $\bm{f}_{n}:=[f(x_{i})]_{i=1}^{n}$, and $\bm{\epsilon}_{n}:=[\epsilon_{i}]_{i=1}^{n}$. The last term on the right-hand side above is sub-Gaussian, since $\mathbb{E}[\bm{\epsilon}_{n}]=0$ and, letting $\bm{\alpha}_{n}:=(\mathbf{K}_{n}+\sigma_{\epsilon}^{2}\mathbf{I})^{-1}\kappa(\pazocal{X}_{n},x)$, we have a sum of independent sub-Gaussian random variables, see e.g. (Pisier, , 2016), Lemma 1.1:
@@
which follows from the definition of sub-Gaussian noise (cf. Assumption B.2). The remaining term on the right-hand side of equation 30 is a zero-mean Gaussian random variable with variance given by:
@@
As equation 30 describes the sum of two independent sub-Gaussian random variables, we can follow similar reasoning to the one applied in equation 31 to show that $\Delta_{n}(x)$ is $s_{n}^{2}(x)$-sub-Gaussian for some $s_{n}^{2}(x)>0$. The resulting sub-Gaussian parameter $s_{n}^{2}(x)$ is then bounded by the sum of the individual sub-Gaussian parameters in equations 31 and 32 as:
@@ B.5 Convergence guarantees
Lemma B.2.
Let assumptions B.1 to B.4 hold, and set $\beta_{t}=\beta>0$, for all $t\in\{0,1,2,\dots\}$. Then the following almost surely holds:
@@
Consider the following upper bound on the predictive variance of a GP model (Steinberg et al., , 2024, Lem. D.3):
@@
where $N_{t}(m,\phi)$ denotes the number of observations collected at $(m,\phi)\in\pazocal{M}\times\Phi$ up to time $t\geq 1$. In addition, letting $\mathfrak{H}_{t}$ denote the $\sigma$-algebra generated by the history of all random variables measurable at time $t$, and setting $\hat{\phi}:=\lim_{t\to\infty}\phi_{t}$, the second Borel-Cantelli lemma (Dubins and Freedman, , 1965) tells us that: [^3] More precisely, the second Borell-Cantelli lemma tells us the two sides are proportional to each other, while equality holds if the right-hand side diverges.
$\forall m\in\pazocal{M},\quad\lim_{t\to\infty}N_{t}(m,\hat{\phi})=\lim_{t\to\infty}\sum_{i=1}^{t}\operatorname*{\mathbb{P}}\left[m_{t}=m\mid\mathfrak{H}_{t-1}\right]\,.$ (37)
Therefore, for $\sigma_{t}^{2}(m,\hat{\phi})\to 0$, we need the series above to diverge. To ensure the latter, we can show that the conditional probabilities in Equation˜39 have a nonzero lower bound or, if they converge to zero, that they do so slowly enough.
@@
where $\lVert f(\cdot,\phi)\rVert_{\infty}=\sup_{m\in\pazocal{M}}|f(m,\phi)|$ denotes the supremum norm of $f(\cdot,\phi)$, and we applied Jensen’s inequality in the last step. Since the kernel $\kappa$ is continuous and bounded, the sub-Gaussian parameter $\sigma_{t}^{2}(\cdot,\phi_{t})$ has a maximum in $\pazocal{M}$, which is finite. As the expected value of the maximum of a finite collection of sub-Gaussian random variables is bounded (see, e.g., Boucheron et al., , 2013, Thr. 2.5), it follows that the GP mean $\mu_{t}$ is almost surely bounded at all times (by, e.g., Markov’s inequality). Considering the model sampling probabilities and that $p_{\min}:=\min_{m\in\pazocal{M}}p(m)>0$ by Assumption B.4, we then have that the following almost surely holds:
@@
Theorem B.3.
Under the assumptions in Lemma˜B.2, we have that the following holds in probability:
$D_{\mathrm{KL}}(q_{u,t}||q_{f,\phi_{t}}^{*})\in\pazocal{O}_{\operatorname*{\mathbb{P}}}(t^{-1/2})\,$, (41)
where for a stochastic process $\{\xi_{t}\}_{t\in{\mathbb{N}}}$, denoting $\xi_{t}\in\pazocal{O}_{\operatorname*{\mathbb{P}}}(g(t))$, for a positive function $g:{\mathbb{N}}\to(0,\infty)$, means that:
@@
Under assumptions B.1 and B.2, given any $\beta>0$, applying standard sub-Gaussian concentration results (Boucheron et al., , 2013) and a union bound, we have that, for all $t\geq 0$:
@@
By Lemma˜B.2, we know that $\sigma_{t}(m,\phi_{t})\in\pazocal{O}(t^{-1/2})$, so that there exists $C>0$ such that $\sigma_{t}(m,\phi_{t})\leq Ct^{-1/2}$, for all $m\in\pazocal{M}$. We then have that:
@@
Appendix C Proof of bijective equivalence between discrete distributions
Proof of Proposition˜3.1.
@@ Appendix C Proof of bijective equivalence between discrete distributions
Appendix D Monte Carlo gradients via score function estimation
@@
By simply choosing $\varsigma=\mathbb{E}_{t\in\{1,\dots,T\}}\left[\pazocal{L}(\phi,\psi)\right]$, where the expectation is estimated online over the iterations of the optimizer, we can reduce variance of $\nabla_{\psi}\pazocal{L}(\phi,\psi)$. See Section˜D.1 for implementation details.
D.1 Control variate for score function estimator
We adopt the approach used in Kingma and Ba, (2017) for obtaining an unbiased running first moment of the loss function. At iteration $t$ we draw $B$ samples $\{m_{t,n}\}_{n=1}^{B}$ and compute
@@ D.1 Control variate for score function estimator
With fixed decay $\beta\in(0,1)$, update the (biased) first moment exactly as in the approach of the Adam optimizer (Kingma and Ba, , 2017):
@@
D.2 Controlling learning rate via the information gain
When using stochastic gradient descent for optimization over parameters of both $q_{\psi}$ and $q_{\phi}$, it is necessary to use careful scaling of the estimated gradients to ensure the optimizer does not “drop off a cliff” into a local minimum. Such phenomena has been observed in related fields such as proximal policy gradients (Schulman et al., , 2017) where the authors demonstrate empirically such a necessity in reinforcement learning problems. In essence, we want to control the learning rate of $\psi$ with respect to the convergence of $\phi\rightarrow\phi^{*}$. We show empirical results for controlling this rate and leave any mathematical properties for the optimal scaling to future research.
@@ D.2 Controlling learning rate via the information gain
When $q_{\psi}$ is a categorical distribution, this quantity is available analytically. However, in general this is not available, but it can be estimated via Monte Carlo integration and importance sampling using available quantities (see Section˜D.3). We choose to set a threshold for the IG between steps, denoted $\beta_{\mathrm{IG}(\psi)}$, and then at each successive step $t$ we scale $\nabla_{\psi}$ using an iterative method such as bisection [^4] In preliminary investigations, other approaches for implementation of this threshold such as constrained optimization and computation of Lagrange multipliers were trialed without success, possibly due to the geometry of the optimization landscape..
D.3 Monte Carlo estimation of information

### v2

@@ Appendix B Analysis of a CoSMIC normalizing flow
(a) $\qquad\tilde{q}_{\phi}^{\triangleleft}(\bm{\theta}_{m},\bm{u}_{\setminus m}\mid m)\;=\;q_{\phi}(\bm{\theta}_{m}\mid m)\;\nu_{d_{\setminus m}}\!\bigl(\bm{u}_{\setminus m}\bigr),$
and so (b) $\qquad\int\tilde{q}_{\phi}^{\triangleleft}(\bm{\theta}_{m},\bm{u}_{\setminus m}\mid m)\,d\bm{u}_{\setminus m}=q_{\phi}(\bm{\theta}_{m}\mid m).$
Write $T_{\phi}^{\triangleleft,-1}=(T_{\phi,m}^{-1},\operatorname{Id})$, where $\operatorname{Id}$ denotes the identity transform, and let the permuted reference vector be $P_{m}\bm{z}=(\bm{z}_{m},\bm{z}_{\setminus m})\in\mathbb{R}^{{d_{m}}}\times\mathbb{R}^{d_{\setminus m}}$. Because the masking function $C$ sets every transform $\tau_{\bm{\rho}_{i}}$ with $C_{i}(m)=0$ to the identity, the inverse flow splits as
$T_{\phi}^{\triangleleft,-1}(\bm{\theta}_{m},\bm{u}_{\setminus m})\;=\;\bigl(T_{\phi,m}^{-1}(\bm{\theta}_{m}),\;\bm{u}_{\setminus m}\bigr),$
where $T_{\phi,m}^{-1}:\Theta_{m}\to\mathbb{R}^{{d_{m}}}$ is the active block and the dummy block is exactly the identity. Consequently, the Jacobian matrix of $T_{\phi}^{\triangleleft,-1}$ is block upper‑triangular with $\det\nabla T_{\phi}^{\triangleleft,-1}=\det\nabla T_{\phi,m}^{-1}\times 1.$
@@
$\tilde{q}_{\phi}^{\triangleleft}(\bm{\theta}_{m},\bm{u}_{\setminus m}\!\mid\!m)$ $=\nu_{{d_{\mathrm{max}}}}\!\bigl(T_{\phi}^{\triangleleft,-1}(\bm{\theta}_{m},\bm{u}_{\setminus m})\bigr)\bigl|\det\nabla T_{\phi}^{\triangleleft,-1}\bigr|$
@@
$=D_{\mathrm{KL}}(q_{\psi,\phi}||\eta)=\pazocal{L}(\psi,\phi).$
Corollary B.1 (Computational complexity).
@@
Appendix C Theoretical analysis of the model weights distribution
@@ C.1 Regularity assumptions
This mild assumption is satisfied, for example, when $\nu$ is a zero-mean Gaussian distribution and $\tilde{f}$ is Lipschitz continuous on its first argument, in which case $\sigma_{\epsilon}$ only depends on $\tilde{f}$ through its Lipschitz constant [Boucheron et al., 2013, Pisier, 2016].
C.2 Gaussian process model
@@
C.3 Upper confidence bound (UCB) algorithm
Given the GP posterior, we formulate an upper confidence bound algorithm [Srinivas et al., 2010] with:
@@ C.3 Upper confidence bound (UCB) algorithm
C.4 Approximation errors under sub-Gaussian noise
@@
where $\kappa(x,\pazocal{X}_{n}):=[\kappa(x,x_{1}),\dots,\kappa(x,x_{n})]$, $\mathbf{K}_{n}:=[\kappa(x_{i},x_{j})]_{i,j=1}^{n}$, $\bm{f}_{n}:=[f(x_{i})]_{i=1}^{n}$, and $\bm{\epsilon}_{n}:=[\epsilon_{i}]_{i=1}^{n}$. The last term on the right-hand side above is sub-Gaussian, since $\mathbb{E}[\bm{\epsilon}_{n}]=0$ and, letting $\bm{\alpha}_{n}:=(\mathbf{K}_{n}+\sigma_{\epsilon}^{2}\mathbf{I})^{-1}\kappa(\pazocal{X}_{n},x)$, we have a sum of independent sub-Gaussian random variables, see e.g. [Pisier, 2016], Lemma 1.1:
@@
which follows from the definition of sub-Gaussian noise (cf. Assumption C.2). The remaining term on the right-hand side of equation 26 is a zero-mean Gaussian random variable with variance given by:
@@
As equation 26 describes the sum of two independent sub-Gaussian random variables, we can follow similar reasoning to the one applied in equation 27 to show that $\Delta_{n}(x)$ is $s_{n}^{2}(x)$-sub-Gaussian for some $s_{n}^{2}(x)>0$. The resulting sub-Gaussian parameter $s_{n}^{2}(x)$ is then bounded by the sum of the individual sub-Gaussian parameters in equations 27 and 28 as:
@@ C.5 Convergence guarantees
Lemma C.2.
Let assumptions C.1 to C.4 hold, and set $\beta_{t}=\beta>0$, for all $t\in\{0,1,2,\dots\}$. Then the following almost surely holds:
@@
Consider the following upper bound on the predictive variance of a GP model [Steinberg et al., 2024, Lem. D.3]:
@@
where $N_{t}(m,\phi)$ denotes the number of observations collected at $(m,\phi)\in\pazocal{M}\times\Phi$ up to time $t\geq 1$. In addition, letting $\mathfrak{H}_{t}$ denote the $\sigma$-algebra generated by the history of all random variables measurable at time $t$, and setting $\hat{\phi}:=\lim_{t\to\infty}\phi_{t}$, the second Borel-Cantelli lemma [Dubins and Freedman, 1965] tells us that: [^4] More precisely, the second Borell-Cantelli lemma shows that the two sides of Equation 33 are proportional to each other, while equality holds if the right-hand side diverges.
$\forall m\in\pazocal{M},\quad\lim_{t\to\infty}N_{t}(m,\hat{\phi})=\lim_{t\to\infty}\sum_{i=1}^{t}\operatorname*{\mathbb{P}}\left[m_{i}=m\mid\mathfrak{H}_{i-1}\right]\,.$ (33)
Therefore, for $\sigma_{t}^{2}(m,\hat{\phi})\to 0$, we need the series above to diverge. To ensure the latter, we can show that the conditional probabilities in Equation˜35 have a nonzero lower bound or, if they converge to zero, that they do so slowly enough.
@@
where $\lVert f(\cdot,\phi)\rVert_{\infty}=\sup_{m\in\pazocal{M}}|f(m,\phi)|$ denotes the supremum norm of $f(\cdot,\phi)$, and we applied Jensen’s inequality in the last step. Since the kernel $\kappa$ is continuous and bounded, the sub-Gaussian parameter $\sigma_{t}^{2}(\cdot,\phi_{t})$ has a maximum in $\pazocal{M}$, which is finite. As the expected value of the maximum of a finite collection of sub-Gaussian random variables is bounded [see, e.g., Boucheron et al., 2013, Thr. 2.5], it follows that the GP mean $\mu_{t}$ is almost surely bounded at all times (by, e.g., Markov’s inequality). Considering the model sampling probabilities and that $p_{\min}:=\min_{m\in\pazocal{M}}p(m)>0$ by Assumption C.4, we then have that the following almost surely holds:
@@
Definition C.1.
Let $\{\xi_{t}\}_{t\in{\mathbb{N}}}$ be a real-valued stochastic process. We say that $\xi_{t}\in\pazocal{O}_{\operatorname*{\mathbb{P}}}(g(t))$, for a positive function $g:{\mathbb{N}}\to(0,\infty)$, if:
@@
Theorem C.3.
Under the assumptions in Lemma˜C.2, we have that the following holds in probability:
$D_{\mathrm{KL}}(q_{u,t}||q_{f,\phi_{t}}^{*})\in\pazocal{O}_{\operatorname*{\mathbb{P}}}(t^{-1/2})\,.$ (39)
@@
Under assumptions C.1 and C.2, given any $\beta>0$, applying standard sub-Gaussian concentration results [Boucheron et al., 2013] and a union bound, we have that, for all $t\geq 0$:
@@
By Lemma˜C.2, we know that $\sigma_{t}(m,\phi_{t})\in\pazocal{O}(t^{-1/2})$, so that there exists $C>0$ such that $\sigma_{t}(m,\phi_{t})\leq Ct^{-1/2}$, for all $m\in\pazocal{M}$. We then have that:
@@
Remark C.4.
The result in Theorem C.3 is similar to Corollary 1 in Oliveira et al. [2021], which also derives the concentration bound for the KL divergence between a surrogate-based approximation of a posterior and the true posterior. However, Oliveira et al.’s result only provides an asymptotic convergence rate requires an upper bound on the information gain of the surrogate model of order $o(\sqrt{t}\;)$ and an appropriately scaled UCB parameter $\beta_{t}$, whereas our result shows that we do not need either of these assumptions whenever a sampling lower bound can be guaranteed, i.e., $\inf_{t\in{\mathbb{N}},m\in\pazocal{M}}\operatorname*{\mathbb{P}}[m_{t+1}=m|\mathfrak{H}_{t}]>0$. In addition, Oliveira et al. [2021] only deals with the static setting where the target posterior does not change over time, while in our case we have a changing $\phi_{t}$ that leads to different targets per optimization step. This non-stationarity requires additional care with the convergence analysis.
Appendix D Bijective equivalence between discrete distributions
Proposition D.1.
Every finite discrete distribution over a finite support $\pazocal{M}=\{m_{1},m_{2},\ldots,m_{k}\}$ has a unique representation as a categorical distribution. Specifically, there exists a bijective mapping between the set of all finite discrete distributions on $\pazocal{M}$ and the set of categorical distributions parameterized by probability vectors $\bm{\psi^{\zeta}}$ over $\pazocal{M}$.
Proof.
@@ Appendix D Bijective equivalence between discrete distributions
Appendix E Monte Carlo gradients via score function estimation
@@
By simply choosing $\varsigma=\mathbb{E}_{t\in\{1,\dots,T\}}\left[\pazocal{L}(\phi,\psi)\right]$, where the expectation is estimated online over the iterations of the optimizer, we can reduce variance of $\nabla_{\psi}\pazocal{L}(\phi,\psi)$. See Appendix E.1 for implementation details.
E.1 Control variate for score function estimator
We adopt the approach used in Kingma and Ba [2017] for obtaining an unbiased running first moment of the loss function. At iteration $t$ we draw $B$ samples $\{m_{t,n}\}_{n=1}^{B}$ and compute
@@ E.1 Control variate for score function estimator
With fixed decay $\beta\in(0,1)$, update the (biased) first moment exactly as in the approach of the Adam optimizer [Kingma and Ba, 2017]:
@@
E.2 Controlling learning rate via the information gain
When using stochastic gradient descent for optimization over parameters of both $q_{\psi}$ and $q_{\phi}$, it is necessary to use careful scaling of the estimated gradients to ensure the optimizer does not “drop off a cliff” into a local minimum. Such phenomena has been observed in related fields such as proximal policy gradients [Schulman et al., 2017] where the authors demonstrate empirically such a necessity in reinforcement learning problems. In essence, we want to control the learning rate of $\psi$ with respect to the convergence of $\phi\rightarrow\phi^{*}$. We show empirical results for controlling this rate and leave any mathematical properties for the optimal scaling to future research.
@@ E.2 Controlling learning rate via the information gain
When $q_{\psi}$ is a categorical distribution, this quantity is available analytically. However, in general this is not available, but it can be estimated via Monte Carlo integration and importance sampling using available quantities (see Appendix E.3). We choose to set a threshold for the IG between steps, denoted $\beta_{\mathrm{IG}(\psi)}$, and then at each successive step $t$ we scale $\nabla_{\psi}$ using an iterative method such as bisection [^5] In preliminary investigations, other approaches for implementation of this threshold such as constrained optimization and computation of Lagrange multipliers were trialed without success, possibly due to the geometry of the optimization landscape..
E.3 Monte Carlo estimation of information